OpenStudy (anonymous):

how do you multiply like variables with exponents? such as, a^5 * a

terenzreignz (terenzreignz):

Same base (variable) and add the exponents :)

OpenStudy (anonymous):

thanks mate :)

terenzreignz (terenzreignz):

\[\Large a^ma^n = a^{m+n}\]

OpenStudy (e.mccormick):

And as a side note: \(\Large (a^m)^n= a^{m\times n}\)